OpenStudy (anonymous):

Given the intersects of a parabola, how do you find p and q in the equation y = a(x-p)(x-q)?

OpenStudy (anonymous):

The X intersects* Meant to specify.

OpenStudy (cwrw238):

p and q are the x intersects a is the value of the y intersect when x = 0

OpenStudy (amistre64):

actually: "apq" would be the value of the y intercept when x=0

OpenStudy (anonymous):

the answer is already given! Because the intercepts(intersects) of the parabola(which is a vertical parabola) y=a(x-p)(x-q) are simply the valuse which go in place of p and q and make the value of y zero! so these values could be found by setting the value of y=0=a(x-p)(x-q) and clearly for this value to be zero either a=0 or x-p=0 or x-q=0 so the solutions are a=0 or x=p or x=q but a=0 implies the value of the y intercept(x=0) and the x interceps(or the points wher the parabola crosses the x axis are) x=p and x=p so the points on the co-ordinate system are (p,0) and (q,0)

OpenStudy (cwrw238):

silly me - thanx amistre :- apq